# for the Hadron Spectrum Collaboration

### v1

@@ IV Scattering Analysis
In coupled-channel scattering with multiple partial waves, the matrices in Eq. (5) are in the space of all hadron-hadron scattering channels and partial waves. There are then multiple elements of the $t$-matrix that are unknown at any value of $E_{cm}$ and must be determined. A convienient parameterization for coupled-channel scattering is the $K$-matrix parameterization
$[t^{-1}(s)]_{\ell SJa,\ell^{\prime}S^{\prime}Jb}=\frac{1}{(2k_{a})^{\ell}}[K^{-1}(s)]_{\ell SJa,\ell^{\prime}S^{\prime}Jb}\frac{1}{(2k_{b})^{\ell^{\prime}}}+\delta_{JJ^{\prime}}\delta_{\ell\ell^{\prime}}\delta_{SS^{\prime}}I_{ab}(s)$ (6)
where the indices run over orbital angular momentum $\ell$, spin $S$, angular momentum $J$ (the $t$-matrix is block diagonal in $J$), and hadron scattering channel $a,b$, with $s=E_{cm}^{2}$ and where the cm-frame momenta $k_{a,b}^{\ell}$ factors promote the expected behavior of partial wave amplitudes close to threshold. Unitarity of the $S$-matrix is ensured if the symmetric matrix $[K^{-1}(s)]_{\ell SJa,\ell^{\prime}S^{\prime}Jb}$ is real for real energies and the diagonal matrix $I_{ab}$ satisfies $\textrm{Im}I_{ab}=-\rho_{a}\delta_{ab}$. The real part of $I_{ab}$ is either chosen to be zero (simple phase space), or a Chew-Mandelstam phase space is used, defined in terms of a dispersive integral as given in App. B of Ref. [59]. There is freedom in how to parameterise the elements of $[K^{-1}(s)]_{\ell SJa,\ell^{\prime}S^{\prime}Jb}$ and here we employ polynominals in $s=E_{cm}^{2}$,
@@ IV.1 ${DD^{\ast}}$ Scattering Below ${D^{\ast}{D}^{\ast}}$ Threshold
The possible partial wave configurations for ${DD^{\ast}}$ are given in Table 1 up to $\ell=3$ and the irreps to which they contribute are given in Tables 7 and 9 in Appendix A, which show partial wave contributions up to $\ell=4$. As we wish to extract the $J^{P}=1^{+}$ scattering amplitudes, we utilize the $[000]T_{1}^{+}$ irrep, where the ${\>\!}^{{3}\!}{S}_{1}$ and ${\>\!}^{{3}\!}{D}_{1,3}$ partial waves contribute. To increase the amount of constraint compared with using $[000]T_{1}^{+}$ alone, we additionally utilize the $[001]$, $[002]$, $[011]$ and $[111]$ $A_{2}$ irreps, which also have contributions from ${\>\!}^{{3}\!}{S}_{1}$ and ${\>\!}^{{3}\!}{D}_{1}$. However, these irreps with overall non zero momentum receive contributions from many other partial waves which must be constrained to extract the $J^{P}=1^{+}$ scattering amplitude, notably the ${\>\!}^{{3}\!}{P}_{0,2}$, ${\>\!}^{{3}\!}{D}_{2,3}$ and ${\>\!}^{{3}\!}{F}_{2,3,4}$ partial waves, which we refer to as background waves [^1] In general, we refer to all partial waves not in $J^{P}=1^{+}$ as background waves.. To this end, we also utilize the $[000]A_{1}^{-}$ and $[000]E^{-}$ irreps, with lowest partial wave contribution ${\>\!}^{{3}\!}{P}_{0}$ and ${\>\!}^{{3}\!}{P}_{2}$, respectively.
@@
The $S$-wave $t$-matrix (in the absence of higher partial waves) can be recast to the form $t=\frac{E_{{cm}}}{2}\frac{1}{k\cot\delta_{0}-ik}$, which makes it obvious that a pole singularity will manifest through an intersection of $k\cot\delta_{0}$ with $ik$ at an energy below threshold. An intersection of $k\cot\delta_{0}$ with $-ik$ corresponds to a pole on the real axis below threshold on the physical sheet, i.e. a bound state. In contrast, an intersection of $k\cot\delta_{0}$ with $ik$ corresponds to a pole below threshold on the unphysical sheet, i.e. a virtual bound state.
@@
Figure 5: (Top) The $S$-wave amplitude from Eq. (8), shown as $(a_{t}k)\cot\delta_{0}$ (green), along with ±$\pm a_{t}|k|$ (blue) for $k^{2}<0$. The inner shaded band reflects the statistical uncertainty while the outer band represents the systematic uncertainty obtained by varying the mass of the $D$ and $D^{\ast}$ mesons and the anisotropy as described in the text. (Bottom) The corresponding phase shift, $\delta_{0}$. Discrete points are determined from individual energy levels using the determinant condition Eq. (5). Those with large statistical error due to their proximity to non interacting levels have been plotted in grey to lessen their visual impact.
@@ IV.2 Coupled-channel ${DD^{\ast}}$, ${D^{\ast}{D}^{\ast}}$ scattering
Building upon the parameterization of elastic ${DD^{\ast}}$ scattering, we now consider higher energies where both the ${DD^{\ast}}$ and ${D^{\ast}{D}^{\ast}}$ channels are kinematically accessible and so we consider coupled-channel ${DD^{\ast}}$-${D^{\ast}{D}^{\ast}}$ scattering. In this higher-energy region, partial waves that are suppressed close to threshold have more phase space and can become important. In order to have sufficient constraint, 109 energy levels are used from both the rest-frame irreps and those with non-zero overall momentum, up to energies of $a_{t}E_{cm}=0.73$. This is far enough below the $DD\pi$ and $D_{0}^{\ast}D$ thresholds to ensure that no three-body effects or contributions from $D_{0}^{\ast}D$ scattering arise.
@@
In this study, we use “singular value limits” in the $\chi^{2}$ minimization. Here the replacement $\lambda_{i}\rightarrow\mathrm{max}(\lambda_{i},\sigma)$ for the eigenvalues, $\lambda_{i}$, of the correlation matrix is made, where $\sigma=\tau\lambda_{1}$, $\tau$ is the singular value limit and $\lambda_{1}$ is the largest eigenvalue of the correlation matrix. This has the effect of lowering the correlations between finite volume energies, but increases the uncertainties [61]. An alternative is to use “singular value cuts”, wherein eigenvalues below $\sigma$ are set to zero, rather than a limit, as was used in Refs. [33, 38]. In practical terms, the approaches are similar, however for a given value of the cutoff $\tau$, “limiting” is found to be less conservative than “cutting”. We choose a value of $\tau=10^{-2}$. For further discussion of the use of singular value limits, see Appendix D.
@@
We must also determine ${D^{\ast}{D}^{\ast}}$ amplitudes in these irreps. There are fewer ${D^{\ast}{D}^{\ast}}$ combinations due to Bose symmetry as discussed in Section II; see Table 2 and the irreps to which they contribute in Tables 8 and 10. In contrast to the ${DD^{\ast}}$ partial waves, we do not include all ${D^{\ast}{D}^{\ast}}$ partial waves with $\ell\leq 2$. Of the allowed ${D^{\ast}{D}^{\ast}}$ partial waves, we consider ${\>\!}^{{3}\!}{S}_{1}$, ${\>\!}^{{3}\!}{D}_{1}$, ${\>\!}^{{5}\!}{P}_{2}$ and ${\>\!}^{{5}\!}{F}_{2}$. Similarly to ${DD^{\ast}}$, the ${D^{\ast}{D}^{\ast}}$ ${\>\!}^{{3}\!}{S}_{1}$ and ${\>\!}^{{3}\!}{D}_{1}$ partial waves contribute to $[000]T_{1}^{+}$ and all non-zero momentum $A_{2}$ irreps, while the ${\>\!}^{{5}\!}{P}_{2}$ and ${\>\!}^{{5}\!}{F}_{2}$ partial waves contribute to $[000]E^{-}$ and all non-zero momentum $A_{2}$ irreps. The choice to include ${\>\!}^{{5}\!}{P}_{2}$ is motivated by the magnitude of the scattering amplitude of ${DD^{\ast}}$ in ${\>\!}^{{3}\!}{P}_{2}$, which was observed to be significant. A similar effect could be expected of the ${D^{\ast}{D}^{\ast}}$ ${\>\!}^{{5}\!}{P}_{2}$ partial wave. Additionally, a fourth amplitude is required to satisfy the quadruple degeneracy in the upper part of the spectrum in the $[111]A_{2}$ irrep (see Figure 2). In Appendix C, we show there is a freedom in the choice of which partial wave to parameterize for, which we choose to be the partial wave ${\>\!}^{{5}\!}{F}_{2}$.
@@
Analogously to the elastic case, we utilize a $K$-matrix parameterization of the form of Eq. (7) to determine simultaneously the $J^{P}=1^{+}$ amplitude and the smaller $J^{P}=0^{-},2^{-},2^{+}$ and $3^{+}$ background waves. Performing this amplitude determination, we find a good description of the spectra using a constant plus linear term in $s$ for the ${DD^{\ast}}$ and ${D^{\ast}{D}^{\ast}}$ $S$-wave amplitudes and constants for all others, [^3] Correlations are present between the parameters in all $J^{P}$. However, for brevity we have chosen to show only the larger diagonal blocks. Parameter correlations between different $J^{P}$ are typically much smaller than those within a given $J^{P}$.
@@
where the first and second uncertainties are as described in Section III.1. We refer to Eq. (9) as the reference parameterization. Very little deviation is observed in the parameters describing the $J^{P}=0^{-},2^{-},2^{+}$ and $3^{+}$ background waves from this determination to when they are determined independently from rest-frame irreps alone. Of particular importance in the reference parameterization is the nonzero value of the matrix element $\gamma^{(0)}_{{DD^{\ast}}\{{\>\!}^{{3}\!}{S}_{1}\}\rightarrow{D^{\ast}{D}^{\ast}}\{{\>\!}^{{3}\!}{S}_{1}\}}$. This matrix element allows the ${DD^{\ast}}\{{\>\!}^{{3}\!}{S}_{1}\}$ and ${D^{\ast}{D}^{\ast}}\{{\>\!}^{{3}\!}{S}_{1}\}$ channels to mix and is critical to obtaining a parameterization suitable for replicating the finite volume spectra, which will be demonstrated in Section V.3.
Figures 7 and 8 display the finite volume spectra obtained from Eq. (5) using the reference parameterization of Eq. (9) for the at rest and in flight irreps, respectively. The lattice energies are well replicated across all irreps. An avoided level crossing occurs in the $[000]T_{1}^{+}$ irrep just below ${D^{\ast}{D}^{\ast}}$ threshold, indicating some coupling between the channels.
@@
Figure 6 (left) shows the amplitudes arising from the reference parameterization of Eq. (9) and for all parameterization variations (right) given in Table 5. In all parameterizations, we observe a rapid opening of the ${DD^{\ast}}$ $S$-wave amplitude at threshold and a pronounced cusp-like feature peaked at ${D^{\ast}{D}^{\ast}}$ threshold. The diagonal ${D^{\ast}{D}^{\ast}}\rightarrow{D^{\ast}{D}^{\ast}}$ and off-diagonal ${DD^{\ast}}\rightarrow{D^{\ast}{D}^{\ast}}$ amplitudes both display a rapid opening at ${D^{\ast}{D}^{\ast}}$ threshold. These three features of the amplitudes suggest the presence of a pole close to the ${D^{\ast}{D}^{\ast}}$ threshold, which would account for the cusp. We also observe that the ${\>\!}^{{3}\!}{D}_{1}$ waves and the coupling between ${\>\!}^{{3}\!}{S}_{1}$ and ${\>\!}^{{3}\!}{D}_{1}$ waves in both hadron-hadron channels are consistent with zero. The background waves in $J^{P}=0^{-},2^{-},2^{+}$ and $3^{+}$ are well constrained in the region near ${D^{\ast}{D}^{\ast}}$ threshold with background waves having a partial wave configuration of $\ell\geq 2$ being statistically consistent with zero. As in the elastic case, we observe that the ${DD^{\ast}}$ amplitudes in the ${\>\!}^{{3}\!}{P}_{0}$ and ${\>\!}^{{3}\!}{P}_{2}$ channels are significantly nonzero, with a large difference in magnitude at higher energies.
To summarize our key findings of these various amplitude variations, as can be seen in Figs. 4 and 6 for both the elastic and coupled-channel cases, we find a strong $S$-wave interaction, moderate $P$-wave interactions and only weak interactions in $D$-wave and higher. The coupled-channel effects are significant between ${DD^{\ast}}\{{\>\!}^{{3}\!}{S}_{1}\}$ and ${D^{\ast}{D}^{\ast}}\{{\>\!}^{{3}\!}{S}_{1}\}$ with a pronounced cusp clearly visible. There is no appreciable coupling found between ${\>\!}^{{3}\!}{S}_{1}$ and ${\>\!}^{{3}\!}{D}_{1}$ in both hadron-hadron channels. The ${DD^{\ast}}\{{\>\!}^{{3}\!}{P}_{0}\}$ is the next most significant amplitude - it shows no sharp features but is large across both elastic and coupled-channel scattering amplitudes. We now turn to investigating the singularities in the scattering amplitudes, from which masses, widths, and channel couplings of any states present can be obtained.

### v2

@@ IV Scattering Analysis
In coupled-channel scattering with multiple partial waves, the matrices in Eq. (5) are in the space of all hadron-hadron scattering channels and partial waves. There are then multiple elements of the $t$-matrix that are unknown at any value of $E_{cm}$ and must be determined. A convenient parameterization for coupled-channel scattering is the $K$-matrix parameterization
$[t^{-1}(s)]_{\ell SJa,\ell^{\prime}S^{\prime}Jb}=\frac{1}{(2k_{a})^{\ell}}[K^{-1}(s)]_{\ell SJa,\ell^{\prime}S^{\prime}Jb}\frac{1}{(2k_{b})^{\ell^{\prime}}}+\delta_{\ell\ell^{\prime}}\delta_{SS^{\prime}}I_{ab}(s)$ (6)
where the indices run over orbital angular momentum $\ell$, spin $S$, angular momentum $J$ (the $t$-matrix is block diagonal in $J$), and hadron scattering channel $a,b$, with $s=E_{cm}^{2}$ and where the cm-frame momenta $k_{a,b}^{\ell}$ factors promote the expected behavior of partial wave amplitudes close to threshold. Unitarity of the $S$-matrix is ensured if the symmetric matrix $[K^{-1}(s)]_{\ell SJa,\ell^{\prime}S^{\prime}Jb}$ is real for real energies and the diagonal matrix $I_{ab}$ satisfies $\textrm{Im}I_{ab}=-\rho_{a}\delta_{ab}$ above threshold. The real part of $I_{ab}$ is either chosen to be zero (simple phase space), or a Chew-Mandelstam phase space is used, defined in terms of a dispersive integral as given in App. B of Ref. [61]. There is freedom in how to parameterise the elements of $[K^{-1}(s)]_{\ell SJa,\ell^{\prime}S^{\prime}Jb}$ and here we employ polynomials in $s=E_{cm}^{2}$,
@@ IV.1 ${DD^{\ast}}$ Scattering Below ${D^{\ast}{D}^{\ast}}$ Threshold
The possible partial wave configurations for ${DD^{\ast}}$ are given in Table 1 up to $\ell=3$ and the irreps to which they contribute are given in Tables 7 and 9 in Appendix A, which show partial wave contributions up to $\ell=4$. As we wish to extract the $J^{P}=1^{+}$ scattering amplitudes, we utilize the $[000]T_{1}^{+}$ irrep, where the ${\>\!}^{{3}\!}{S}_{1}$ and ${\>\!}^{{3}\!}{D}_{1,3}$ partial waves contribute. To increase the amount of constraint compared with using $[000]T_{1}^{+}$ alone, we additionally utilize the $[001]$, $[002]$, $[011]$ and $[111]$ $A_{2}$ irreps, which also have contributions from ${\>\!}^{{3}\!}{S}_{1}$ and ${\>\!}^{{3}\!}{D}_{1}$. However, these irreps with overall non zero momentum receive contributions from many other partial waves which must be constrained to extract the $J^{P}=1^{+}$ scattering amplitude, notably the ${\>\!}^{{3}\!}{P}_{0,2}$, ${\>\!}^{{3}\!}{D}_{2,3}$ and ${\>\!}^{{3}\!}{F}_{2,3,4}$ partial waves. To this end, we also utilize the $[000]A_{1}^{-}$ and $[000]E^{-}$ irreps, with lowest partial wave contribution ${\>\!}^{{3}\!}{P}_{0}$ and ${\>\!}^{{3}\!}{P}_{2}$, respectively.
@@
The $S$-wave $t$-matrix (in the absence of higher partial waves) can be recast to the form $t=\frac{E_{{cm}}}{2}\frac{1}{k\cot\delta_{0}-ik}$, which makes it obvious that a pole singularity will manifest through an intersection of $k\cot\delta_{0}$ with ±$\pm ik$ at an energy below threshold. An intersection of $k\cot\delta_{0}$ with $-ik$ corresponds to a pole on the real axis below threshold on the physical sheet, i.e. a bound state. In contrast, an intersection of $k\cot\delta_{0}$ with $ik$ corresponds to a pole below threshold on the unphysical sheet, i.e. a virtual bound state.
@@
Figure 5: (Top) The $S$-wave amplitude from Eq. (8), shown as $(a_{t}k)\cot\delta_{0}$ (green), along with ±$\pm a_{t}|k|$ (blue) for $k^{2}<0$. The inner shaded band reflects the statistical uncertainty while the outer band represents the systematic uncertainty obtained by varying the mass of the $D$ and $D^{\ast}$ mesons and the anisotropy as described in the text. (Bottom) The corresponding phase shift, $\delta_{0}$. Discrete points are determined from individual energy levels using the determinant condition Eq. (5). Those with large statistical error due to their proximity to non interacting levels have been plotted in grey to lessen their visual impact. The vertical dashed red line in both plots denotes the position of the beginning of the left hand cut due to one pion exchange between the $D$ and $D^{\ast}$.
@@ IV.2 Coupled-channel ${DD^{\ast}}$, ${D^{\ast}{D}^{\ast}}$ scattering
Building upon the parameterization of elastic ${DD^{\ast}}$ scattering, we now consider higher energies where both the ${DD^{\ast}}$ and ${D^{\ast}{D}^{\ast}}$ channels are kinematically accessible and so we consider coupled-channel ${DD^{\ast}}$-${D^{\ast}{D}^{\ast}}$ scattering. In this higher-energy region, partial waves that are suppressed close to threshold have more phase space and can become important. In order to have sufficient constraint, 109 energy levels are used from both the rest-frame irreps and those with non-zero overall momentum, up to energies of $a_{t}E_{cm}=0.73$. This is far enough below the $DD\pi$ threshold to ensure that no three-body effects arise [^3] This also ensures that no effects from $D_{0}^{\ast}D$ scattering contribute, as the $D_{0}^{\ast}D$ threshold opens just below the $DD\pi$ threshold..
@@
In this study, we use “singular value limits” in the $\chi^{2}$ minimization. Here the replacement $\lambda_{i}\rightarrow\mathrm{max}(\lambda_{i},\sigma)$ for the eigenvalues, $\lambda_{i}$, of the correlation matrix is made, where $\sigma=\tau\lambda_{1}$, $\tau$ is the singular value limit and $\lambda_{1}$ is the largest eigenvalue of the correlation matrix. This has the effect of lowering the correlations between finite volume energies, but increases the uncertainties [63]. An alternative is to use “singular value cuts”, wherein modes corresponding to eigenvalues of the correlation matrix below $\sigma$ are removed, this was used in Refs. [33, 38]. In practical terms, the approaches are similar, however for a given value of the cutoff $\tau$, “limiting” is found to be less conservative than “cutting”. We choose a value of $\tau=10^{-2}$. For further discussion of the use of singular value limits, see Appendix D.
@@
We must also determine ${D^{\ast}{D}^{\ast}}$ amplitudes in these irreps. There are fewer ${D^{\ast}{D}^{\ast}}$ combinations due to Bose symmetry as discussed in Section II; see Table 2 and the irreps to which they contribute in Tables 8 and 10. In contrast to the ${DD^{\ast}}$ partial waves, we do not include all ${D^{\ast}{D}^{\ast}}$ partial waves with $\ell\leq 2$. Of the allowed ${D^{\ast}{D}^{\ast}}$ partial waves, we consider ${\>\!}^{{3}\!}{S}_{1}$, ${\>\!}^{{3}\!}{D}_{1}$, ${\>\!}^{{5}\!}{P}_{2}$ and ${\>\!}^{{5}\!}{F}_{2}$. Similarly to ${DD^{\ast}}$, the ${D^{\ast}{D}^{\ast}}$ ${\>\!}^{{3}\!}{S}_{1}$ and ${\>\!}^{{3}\!}{D}_{1}$ partial waves contribute to $[000]T_{1}^{+}$ and all non-zero momentum $A_{2}$ irreps, while the ${\>\!}^{{5}\!}{P}_{2}$ and ${\>\!}^{{5}\!}{F}_{2}$ partial waves contribute to $[000]E^{-}$ and all non-zero momentum $A_{2}$ irreps. The choice to include ${\>\!}^{{5}\!}{P}_{2}$ is motivated by the magnitude of the scattering amplitude of ${DD^{\ast}}$ in ${\>\!}^{{3}\!}{P}_{2}$, which was observed to be significant in the elastic analysis of Section IV.1 (see Figure 4). A similar effect could be expected of the ${D^{\ast}{D}^{\ast}}$ ${\>\!}^{{5}\!}{P}_{2}$ partial wave. Additionally, a fourth amplitude is required to satisfy the quadruple degeneracy in the upper part of the spectrum in the $[111]A_{2}$ irrep (see Figure 2). In Appendix C, we show there is a freedom in the choice of which partial wave to parameterize for, which we choose to be the partial wave ${\>\!}^{{5}\!}{F}_{2}$.
@@
Analogously to the elastic case, we utilize a $K$-matrix parameterization of the form of Eq. (7) to determine simultaneously the $J^{P}=1^{+}$ amplitude and the smaller partial waves in $J^{P}=0^{-},2^{-},2^{+}$ and $3^{+}$. Performing this amplitude determination, we find a good description of the spectra using a constant plus linear term in $s$ for the ${DD^{\ast}}$ and ${D^{\ast}{D}^{\ast}}$ $S$-wave amplitudes and constants for all others, [^4] Correlations are present between the parameters in all $J^{P}$. However, for brevity we have chosen to show only the larger diagonal blocks. Parameter correlations between different $J^{P}$ are typically much smaller than those within a given $J^{P}$.
@@
where the first and second uncertainties are as described in Section IV.1. We refer to Eq. (9) as the reference parameterization. Very little deviation is observed in the parameters describing the partial waves in $J^{P}=0^{-},2^{-},2^{+}$ and $3^{+}$ from this determination to when they are determined independently from rest-frame irreps alone. Of particular importance in the reference parameterization is the nonzero value of the matrix element $\gamma^{(0)}_{{DD^{\ast}}\{{\>\!}^{{3}\!}{S}_{1}\}\rightarrow{D^{\ast}{D}^{\ast}}\{{\>\!}^{{3}\!}{S}_{1}\}}$. This matrix element allows the ${DD^{\ast}}\{{\>\!}^{{3}\!}{S}_{1}\}$ and ${D^{\ast}{D}^{\ast}}\{{\>\!}^{{3}\!}{S}_{1}\}$ channels to mix and is critical to obtaining a parameterization suitable for replicating the finite volume spectra, which will be demonstrated in Section V.3.
Figures 6 and 7 display the finite volume spectra obtained from Eq. (5) using the reference parameterization of Eq. (9) for the at rest and in flight irreps, respectively. The lattice energies are well replicated across all irreps. An avoided level crossing occurs in the $[000]T_{1}^{+}$ irrep just below ${D^{\ast}{D}^{\ast}}$ threshold, indicating some coupling between the channels.
@@
Figure 8 (left) shows the amplitudes arising from the reference parameterization of Eq. (9) and for all parameterization variations (right) given in Table 5. In all parameterizations, we observe a rapid opening of the ${DD^{\ast}}$ $S$-wave amplitude at threshold and a pronounced cusp-like feature peaked at ${D^{\ast}{D}^{\ast}}$ threshold. The diagonal ${D^{\ast}{D}^{\ast}}\rightarrow{D^{\ast}{D}^{\ast}}$ and off-diagonal ${DD^{\ast}}\rightarrow{D^{\ast}{D}^{\ast}}$ amplitudes both display a rapid opening at ${D^{\ast}{D}^{\ast}}$ threshold. These three features of the amplitudes suggest the presence of a pole close to the ${D^{\ast}{D}^{\ast}}$ threshold, which would account for the cusp. We also observe that the ${\>\!}^{{3}\!}{D}_{1}$ waves and the coupling between ${\>\!}^{{3}\!}{S}_{1}$ and ${\>\!}^{{3}\!}{D}_{1}$ waves in both hadron-hadron channels are consistent with zero. The other waves in $J^{P}=0^{-},2^{-},2^{+}$ and $3^{+}$ are well constrained in the region near ${D^{\ast}{D}^{\ast}}$ threshold with other waves having a partial wave configuration of $\ell\geq 2$ being statistically consistent with zero. As in the elastic case, we observe that the ${DD^{\ast}}$ amplitudes in the ${\>\!}^{{3}\!}{P}_{0}$ and ${\>\!}^{{3}\!}{P}_{2}$ channels are significantly nonzero, with a large difference in magnitude at higher energies.
To summarize our key findings of these various amplitude variations, as can be seen in Figs. 4 and 8 for both the elastic and coupled-channel cases, we find a strong $S$-wave interaction, moderate $P$-wave interactions and only weak interactions in $D$-wave and higher. The coupled-channel effects are significant between ${DD^{\ast}}\{{\>\!}^{{3}\!}{S}_{1}\}$ and ${D^{\ast}{D}^{\ast}}\{{\>\!}^{{3}\!}{S}_{1}\}$ with a pronounced cusp clearly visible. There is no appreciable coupling found between ${\>\!}^{{3}\!}{S}_{1}$ and ${\>\!}^{{3}\!}{D}_{1}$ in both hadron-hadron channels. The ${DD^{\ast}}\{{\>\!}^{{3}\!}{P}_{0}\}$ is the next most significant amplitude - it shows no sharp features but is large across both elastic and coupled-channel scattering amplitudes. We now turn to investigating the singularities in the scattering amplitudes, from which masses, widths, and channel couplings of any states present can be obtained.